### v1

@@ 2.1.1 Slow migration
When the planets undergo fast convergent migration, they directly bypass the resonance before the eccentricities reach to the equilibrium values. In this case $\dot{a}/a$ is driven by $\tau_{a}$ but not $\tau_{e}$, and Eq. (3) can be rewritten as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{2\tau_{a}},$ (6)
where the e-damping term is ignored. Here the eccentricities are purely excited by the resonant term, which can be deduced from the classical pendulum model (see appendix Sec. Appendix A: The derivation of the excited eccentricity in the classical pendulum model for derivation)
@@
Combing Eqs. (6) and (7), the resonance trapping criterion of slow migration can be obtained as
$\tau_{a}>\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}.$ (8)
@@
$\begin{cases}{\displaystyle\tau_{a}>\frac{1}{2n_{\rm i}(3(j-1))^{1/3}}\frac{1}{(\mu_{\rm o}\alpha f_{\rm d,i})^{4/3}}\ \sim\mu_{\rm o}^{-4/3}}\ \ \hfill\mbox{ at $q\ll 1$},\\ {\displaystyle\tau_{a}>\frac{1}{2n_{\rm o}(3j)^{1/3}}\frac{1}{(\mu_{\rm i}f_{\rm d,o}^{{}^{\prime}})^{4/3}}\ \sim\mu_{\rm i}^{-4/3}}\ \ \ \hfill\mbox{at $q\gg 1$}.\\ \end{cases}$ (9)
This indicates that the threshold migration rate increases with the planet mass. Intuitively, the planet resonant interaction becomes stronger as their masses increase. Therefore, the planets require faster migration in order to break the resonance.
@@ 2.1.2 Weak eccentricity damping
When the planets migrate relatively slow, they have time to reach the equilibrium state. Substituting the equilibrium eccentricity into the right side of Eq. (3), it turns out that the magnitude of e-damping term is comparable to that of the a-decay term. In this circumstance, the e-damping term becomes a key factor for resonance trapping.
Considering the limit of $\sin{\varphi_{\rm i}}{=}1$ and $\sin{\varphi_{\rm o}}{=}-1$, Eq. (3) can be rewritten as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$ (10)
@@
where the left-hand side of the planetary interaction maintains a finite magnitude at resonances (${\propto}\mu ne$). If eccentricity damping is very strong, the orbital circularization leads large $\dot{a}/a$,, and the resonance can no longer hold. One example of resonance crossing is shown in Figure 1. Hence, resonance trapping requires weak e-damping. Substituting Eq. (10) with equilibrium values of $e_{\rm i}$ and $e_{\rm o}$ from Eqs. (4a) and (4b), we obtain the resonance trapping criterion in the weak e-damping regime as
$\tau_{e}\tau_{a}>\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}.$ (11)
In the limit of $q{\ll}1$, the trapping criterion can be simplified as
$\tau_{e}\tau_{a}>\frac{1}{2j(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i})^{2}},$ (12)
which is in consistent with Eq. (25) of Huang & Ormel (2023).
@@
Altogether, the resonance trapping requires both relatively weak eccentricity damping and slow migration. Taking Eqs. (11) and (8) together, we formulate the necessary condition for $\tau_{a}$ and $\tau_{e}$ as
@@
$\begin{cases}{\displaystyle(\tau_{e}\tau_{a})_{\rm weak,e}=\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vspace{0.1cm}\\ {\displaystyle(\tau_{a})_{\rm slow,mig}=\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}].\end{cases}$ (13)
Figure 3: Resonance crossing of a two-planet system with convergent migration. The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, and the migration parameters are $\tau_{a}=4\times 10^{5}yr$ and $\tau_{a}/\tau_{e}=5\times 10^{2}$. The planets are trapped in the $2:1$ resonance at $t{\simeq}3\times 10^{4}$ yr and their eccentricities first excite to $2\times 10^{-2}$, and starts to exhibit significant oscillations around its equilibrium value. The resonant angle of the inner planet remains at the equilibrium value and begins to oscillate after some time. In phase space, the inner planet initially evolves to a non-zero eccentricity resonant fixed point and subsequently starts oscillating around it. Over time, the system settles into a limit cycle around the fixed point, reflecting the dynamic equilibrium state of the resonance.
@@ 2.2 Post-evolution after resonant trapping
Here we assume that the planet pairs have satisfied the resonance trapping conditions in Sec. LABEL:sec:riteria and study their following dynamical evolution. The goal is to explore under which configuration the resonance state is temporary or long-term stable.
Our approach follows the analysis outlined in Section 2 of Xu & Lai (2017) (also see Chapter 8.8 of Murray & Dermott 1999). We first focus on the situation that the mass of the inner planet is much lower than the outer one ($m_{\rm i}{\ll}m_{\rm o}$, $q{\ll}1$). In the absence of planet-disk interactions, the system’s Hamiltonian $H$ near the $j$:$j-1$ MMR can be expressed as
$-H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (14)
@@
$\Phi$ $=$ $[\frac{3(j-1)^{2}}{f_{\rm d,i}\mu_{\rm o}\alpha_{\rm 0}}]^{2/3}\frac{1-\sqrt{1-e_{\rm i}^{2}}}{2}\sim\mu_{\rm o}^{2/3}e_{\rm i}^{2},$ (15b)
$\eta$ $=$ $[\frac{1}{\sqrt{3}(j-1)f_{\rm d,i}\mu_{\rm o}}]^{2/3}[(j-1)\alpha_{\rm 0}^{-2/3}-j\alpha_{\rm 0}^{5/6}].$ (15c)
The first two parameters are a pair of conjugate coordinate and momentum, where $\varphi$ is the resonant angle and $\Phi$ represents the angular momentum deficit (AMD). The third parameter, $\eta$ is the dimensionless that quantifies the deviation level from the exact resonance. The orbital period ratio of the two planets is the exact integer ratio at $\eta{=}0$. The initial semi-major axis ratio of the two planet $\alpha_{\rm 0}\equiv\alpha[1+(j-1)e_{\rm i}^{2}]$ is the conjugate momentum of a fast angle, which can be derived from Eq.8-92 of Murray & Dermott (1999). We note that in the absence of planet-disk interactions, the parameters $\eta$ and $\alpha_{0}$ remain constant.
As we will demonstrate later, the effects of planet-disk interaction ($\tau_{\rm a},\tau_{\rm e}$) determines the stability of the planet pair in resonance.
@@
$\dot{\Phi}{=}-\frac{\partial H}{\partial\varphi}+\frac{\partial\Phi}{\partial e}\frac{de}{dt},\ \ \ \dot{\varphi}{=}\frac{\partial H}{\partial\Phi}.$ (16)
By incorporating the orbital decay and eccentricity damping terms in Eq.(1), Eq. (16) can be derived as
@@
$\dot{\eta}$ $=$ $\frac{\partial\eta}{\partial\alpha_{0}}\frac{d\alpha_{0}}{dt}{=}-\frac{\beta}{\tau_{a}}+\frac{2(p+2j-2)\Phi}{(j-1)\tau_{e}},$ (17c)
@@
We then use the linear perturbation method to calculate these quantities ($\Phi{=}\Phi_{\rm eq}+\delta\Phi$ etc.). The corresponding linearized equations read
${\begin{bmatrix}\delta\dot{\Phi}\\ \delta\dot{\varphi}\\ \delta\dot{\eta}\end{bmatrix}=\begin{bmatrix}-\tau_{e}^{-1}&\Phi_{\rm eq}^{1/2}&0\\ -(0.25\Phi^{-3/2}+2)&-\tau_{e}^{-1}&-1\\ 4j\tau_{\rm e}^{-1}(j-1)^{-1}&0&0\end{bmatrix}\begin{bmatrix}\delta{\Phi}\\ \delta{\varphi}\\ \delta{\eta}\end{bmatrix}}.$ (20)
@@
$\lambda^{3}+\frac{2}{\tau_{e}}\lambda^{2}+\left(2\Phi_{\rm eq}^{1/2}+\frac{1}{4\Phi_{\rm eq}}\right)\lambda+\frac{\beta}{\tau_{a}\Phi_{\rm eq}^{1/2}}=0.$ (21)
We provide a general approach to solve the real part solutions of Eq. (21) as folows. Suppose that a cubic equation has a form of $A\lambda^{3}+B\lambda^{2}+C\lambda+D{=}0$ where $A,B,C,D{>}0$. Assuming that $\lambda{=}\lambda_{\rm r}\pm i\lambda_{\rm i}$, it can be proved that when $D{<}BC/A$, the real part of eigenvalues $\lambda_{r}{<}0$. This consequently means that the trapping is stable where the trajectory of Hamiltonian $H$ converges to the fixed point (see an illustration in Figure 2). Substituting $A,B,C,D$ with Eq. (21), we obtain our stability criterion as
$\frac{\tau_{a}}{\tau_{e}}>\left(\frac{3}{f_{\rm d,i}}\right)^{2/3}\frac{(j-1)^{2/9}j^{4/9}}{2j}\mu_{\rm o}^{-2/3}\sim\mu_{\rm o}^{-2/3}.$ (22)
@@
Our analysis can also be extended to the case when the mass of the outer planet is much lower than that of the inner one ($m_{\rm i}{\gg}m_{\rm o}$, $q{\gg}1$). In such a circumstance, we find that the real part of eigenvalues of characteristic equation $\lambda_{\rm r}{<}0$ always holds. Importantly, this means that once the planet pair gets trapped in the resonance, the trapping is always stable as long as $m_{\rm i}{\gg}m_{\rm o}$.
Figure 4: Level curves of the Hamiltonian for different $\eta$, plotted in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$. The black (red) dots mark the stable (unstable) fixed points, the thick black line in panel (b) marks the separatrix. Each curve represents the conserved Hamiltonian in the absence of dissipitive disk forces. The color of the lines represents the values of the Hamiltonian. When considering the planet-disk interaction, during the convergent migration of the planet, H increases, which corresponds to a color shift from purple and green to yellow. When overstability occurs, the perturbations grow, causing the curve to expand, and H decreases, corresponding to a color shift from yellow and green to purple.

### v2

@@ 2.1.1 Slow migration
Relatively slow migration is a necessary resonant capture condition. When the planets undergo fast convergent migration, they directly bypass the resonance before the eccentricities reach their equilibrium values. In this case $\dot{a}/a$ is driven by $\tau_{\rm m}$ but not $\tau_{e}$, and Eq. (3) can be rewritten as
$(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{\tau_{\rm m}},$ (6)
where the e-damping term is ignored. We adopt $\sin{\varphi_{\rm i}}{=}-\sin{\varphi_{\rm o}}{=}1$, so the left-hand side of Eq. 6 represents the maximum strength of the resonant interaction. From the classical pendulum model (see Appendix A for the derivation), the maximum value of $e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}$ can be obtained as
@@
If the migration rate is too rapid, the migration term in Eq. 6 always exceeds the resonant term, and therefore the equilibrium state cannot be reached. Combing Eqs. (6) and (7), the resonance trapping criterion of slow migration can be written as
$\tau_{\rm m}>\frac{1}{[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}.$ (8)
@@
[…]
This indicates that the threshold migration rate increases with the planet mass. Intuitively, the planet’s resonant interaction becomes stronger as their masses increase. Therefore, the planets require faster migration in order to break the resonance.
@@ 2.1.2 Weak eccentricity damping
When the planets migrate relatively slowly towards each other, they have time to reach the equilibrium state. Substituting the equilibrium eccentricity into the right side of Eq. (3), it turns out that the magnitude of the e-damping term is comparable to that of the angular momentum damping term. In this circumstance, the e-damping term becomes a key factor for resonance trapping.
By substituting $e_{\rm i,eq}$ and $e_{\rm o,eq}$ into Eqs.(2c) and (2d), $\sin{\varphi_{\rm i,o}}$ can be calculated. Naturally, the absolute values of $\sin{\varphi_{\rm i,o}}$ cannot exceed $1$. Therefore, by considering the limit of $\sin{\varphi_{\rm i}}{=}1$ and $\sin{\varphi_{\rm o}}{=}-1$, Eq.(3) can be rewritten as
$(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{\tau_{\rm m}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}}.$ (10)
@@
If eccentricity damping is very strong, the orbital circularization leads to large $\dot{a}/a$, and the resonance can no longer hold. One example of resonance crossing is shown in Figure 2. Hence, resonance trapping requires weak e-damping. Substituting Eq. (10) with equilibrium values of $e_{\rm i}$ and $e_{\rm o}$ from Eqs. (4a) and (4b), we obtain the resonance trapping criterion in the weak e-damping regime as
$\tau_{e}\tau_{\rm m}>\frac{1}{(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}.$ (11)
The expression agrees with the one obtained by Batygin & Petit (2023), who used a different yet equivalent analytical method. In the limit of $q{\ll}1$, the trapping criterion can be simplified as
$\tau_{e}\tau_{\rm m}>\frac{1}{j(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i})^{2}},$ (12)
consistent with Eq. (25) of Huang & Ormel (2023).
@@
Altogether, resonance trapping requires both relatively weak eccentricity damping and slow migration. Taking Eqs. (11) and (8) together, we formulate the threshold condition for $\tau_{\rm m}$ and $\tau_{e}$ as
@@
$\begin{cases}{\displaystyle\tau_{e}\tau_{\rm m}>\frac{1}{(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vskip 2.84544pt\\ {\displaystyle\tau_{\rm m}>\frac{1}{[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}].\end{cases}$ (13)
Figure 3: Overstable resonance capture of a two-planet system with convergent migration. The planet and disk parameters are $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $\tau_{\rm m}{=}8\times 10^{5}$ yr and $\tau_{\rm m}/\tau_{e}{=}10^{3}$. By $t{\simeq}3\times 10^{4}$ yr, the planets become trapped in a 22:11 MMR. The inner planet’s eccentricity rapidly excites to $0.02$, followed by oscillations with significant amplitude, while the resonant angle exhibits similar behavior. In phase space, the trajectory initially evolves toward a resonant equilibrium point (black dot) before stabilizing into a limit cycle around this point. We term this phenomenon-where the system remains trapped by librating around an equilibrium with the finite amplitude-an overstable trap. See https://github.com/llh-astro/resonance/raw/main/gif/overstable_trap.gif
@@ 2.2 Post-evolution after resonant trapping
Here we assume that the planet pairs have satisfied the resonance trapping conditions in Section 2.1 and study their following dynamical evolution. The goal is to explore under which configuration the resonance state is temporary or long-term stable.
Our approach follows the analysis outlined in Section 2 of Xu & Lai (2017) (also see Chapter 8.8 of Murray & Dermott 1999). We first focus on the situation that the mass of the inner planet is much lower than the outer one ($m_{\rm i}{\ll}m_{\rm o}$, $q{\ll}1$). In the absence of planet-disk interactions, the system’s Hamiltonian $H$ near the $j$:j1j-1 MMR can be expressed as
$H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (14)
@@
$\Phi$ $=$ $\left[\frac{3(j-1)^{2}}{f_{\rm d,i}\mu_{\rm o}\alpha_{\rm 0}}\right]^{2/3}\frac{1-\sqrt{1-e_{\rm i}^{2}}}{2}\sim\mu_{\rm o}^{2/3}e_{\rm i}^{2},$ (15b)
$\eta$ $=$ $\left[\frac{1}{\sqrt{3}(j-1)f_{\rm d,i}\mu_{\rm o}}\right]^{2/3}[(j-1)\alpha_{\rm 0}^{-2/3}-j\alpha_{\rm 0}^{5/6}].$ (15c)
The first two parameters are a pair of conjugate coordinates and momentum, where $\varphi$ is the resonant angle and $\Phi$ represents the angular momentum deficit (AMD). The third parameter, $\eta$, is dimensionless and quantifies the deviation level from the exact resonance. The orbital period ratio of the two planets is the exact integer ratio at $\eta{=}0$. The initial semi-major axis ratio of the two planets $\alpha_{\rm 0}\equiv\alpha[1+(j-1)e_{\rm i}^{2}]$ is the conjugate momentum of a fast angle, which can be derived from Eq.8-92 of Murray & Dermott (1999). We note that in the absence of planet-disk interactions, the parameters $\eta$ and $\alpha_{0}$ remain constant.
@@
$\dot{\Phi}{=}-\frac{\partial H}{\partial\varphi}+\frac{\partial\Phi}{\partial e}\frac{de}{dt}+\frac{\partial\Phi}{\partial a}\frac{da}{dt},\ \ \ \dot{\varphi}{=}\frac{\partial H}{\partial\Phi}.$ (16)
Since $\Phi$ is independent of $a$ ($\alpha_{0}$ is a constant), the term involving $\partial\Phi/\partial a=0$ can be neglected. By incorporating the angular momentum damping and eccentricity damping terms in Eq.(1), Eq. (16) can be derived as
@@
$\dot{\eta}$ $=$ $\frac{\partial\eta}{\partial\alpha_{0}}\frac{d\alpha_{0}}{dt}{=}-\frac{2\beta}{\tau_{\rm m}}+\frac{4j\Phi}{(j-1)\tau_{e}},$ (17c)
@@
We then use the linear perturbation method to analyze the system’s dynamics ($\Phi{=}\Phi_{\rm eq}+\delta\Phi$ etc.), and the corresponding linearized equations read
[…]
@@
$\lambda^{3}+\frac{2}{\tau_{e}}\lambda^{2}+\left(2\Phi_{\rm eq}^{1/2}+\frac{1}{4\Phi_{\rm eq}}\right)\lambda+\frac{2\beta}{\tau_{\rm m}\Phi_{\rm eq}^{1/2}}=0.$ (21)
We provide a general approach to solve the real part solutions of Eq. (21) as follows. Suppose that a cubic equation has a form of $A\lambda^{3}+B\lambda^{2}+C\lambda+D{=}0$ where A,B,C,D>0A,B,C,D{>}0. Assuming that $\lambda{=}\lambda_{\rm r}\pm i\lambda_{\rm i}$, it can be proved that when D<BC/AD{<}BC/A, the real part of eigenvalues $\lambda_{r}{<}0$. This consequently means that the trapping is stable where the trajectory of Hamiltonian $H$ converges to the equilibrium (see an illustration in Figure 2). Substituting $A,B,C,D$ with Eq. (21), we obtain the stability criterion as
$\frac{\tau_{\rm m}}{\tau_{e}}>\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}\frac{1}{j}\left[\frac{(j-1)}{f_{\rm d,i}\alpha}\right]^{2/3}\sim\mu_{\rm o}^{-2/3}.$ (22)
@@
Our analysis can also be extended to the case when the mass of the outer planet is much lower than that of the inner one ($m_{\rm i}{\gg}m_{\rm o}$, $q{\gg}1$). In such a circumstance, we find that the real part of eigenvalues of the characteristic equation $\lambda_{\rm r}{<}0$ always holds. Importantly, this means that once the planet pair gets trapped in the resonance, the trapping is always stable as long as $m_{\rm i}{\gg}m_{\rm o}$. This holds for the planet pairs with an equal eccentricity damping timescale. Detailed analysis for the planet pairs with different $\tau_{\rm e}$ can be found in Deck & Batygin (2015).
Figure 4: Level curves of the Hamiltonian (Eq. 14) for different $\eta$, plotted in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$. The black (red) dots mark the stable (unstable) fixed points, the thick black line in panel (b) marks the separatrix. Each curve represents the conserved Hamiltonian in the absence of dissipitive disk forces. The color of the lines represents the values of the Hamiltonian.